a particle is thrown vertically upwards with velocity such that time of flight is more than 2 sec. Prove that in the last second of its upward journey it will travel a distance 5m.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
T = 2 = 2u/g => u = 9.8 m/s (minimum velocity)

in last second .i.e T -1 = 2-1 = 1 second,

h1 = 9.8*2 -0.5*g*4
h2 = 9.8*1 – 0.5*g*1
Answer is h2 – h1
rest u can do!